### v2

# A Dark Matter Fermionic Quantum Fluid from Standard Model Dynamics

Abstract
We present a model of dark matter as a superconducting fluid of Cooper pairs of right handed neutrinos or of vector-like quarks. The superconducting dark matter is induced by attractive channels in the Standard Model Higgs and color sectors of the Standard Model, respectively. We show that, for each case, the solution to the gap equation provides viable dark matter candidates for suitable chemical potential values. The mechanism yields an ultra-light neutrino condensate with a mass of $m_{\rm DM}\sim 10^{-19}\text{eV}$ or a vector-like quark condensate with a wide range of possible masses. Both cosmological and particle physics constraints on the model lead to a connection between the number of effective relativistic species $N_{\rm eff}$, and the chemical potential and CMB temperature at the time of fermion creation. For certain values of the fermionic chemical potential, we find a relation between the superconducting fermion and baryon densities, with implications for the coincidence between the dark matter and baryon densities in standard cosmology. Given the natural eV scale of neutrinos, this mechanism may have implications for the Hubble tension.
preprint: APS/123-QED
I Introduction
The existence of dark matter (DM) is now a settled component of the concordance model of cosmology and concrete evidence of physics beyond the Standard Model (SM) [1]. Observations ranging from galactic rotation curves, large scale structure, to the cosmic microwave background (CMB) radiation have established evidence both for the existence of DM and its impact on visible matter, and on the cosmological history of the universe [2]. However, despite the robust evidence for its existence, the microphysical nature and identity DM are still unknown.
In this paper we propose a minimal model of dark matter which relies on the universality of the pairing mechanism of Bardeen-Cooper-Schrieffer (BCS) theory (Cooper pairs) induced by attractive interations given by either the Higgs or QCD sector of the Standard Model. The idea that the dark matter can be a BCS condensate was first discussed by the authors [3, 4, 5], and has been explored in other cosmological contexts [6], however its connection to the Standard Model was unexplored. In this manner, the dark matter is a stable massive Higgs mode of a superconducting state of fermions. As we will see, the resulting dynamics predicts a new relationship between the concordance of the CMB temperature, $T_{\rm CMB}$ and the chemical potential of the fermions (at the time of their creation) which form the dark matter superconductor.
Our description of the dark matter will have implications for extensions, if any, of the Standard Model and could give indications for further experimental detections to discriminate amongst various candidate models. Importantly, any viable microphysical description must satisfy a few criteria. (1) It must be dark, with a small a cross-section-to-mass ratio $\tfrac{\sigma}{M}\ll 1$ for SM-DM processes. It is imperative that the microscopic model explain why ordinary SM interactions are absent or small. (2) The condensate must satisfy the observational constraints on dark matter, such as a small self-interaction rate and a dust-like equation of state.
Beyond these basic criteria, a successful microphysical description must further explain observational features. Both WMAP and more recently, Planck have measured a rough parity between the dark matter energy density $\Omega_{c}$ and baryonic energy density $\Omega_{b}$, namely $\Omega_{c}\sim 5\Omega_{b}$ [7, 8]. Classes of models exist which seek to explain this apparent coincidence. One such class is known as asymmetric dark matter (ADM) [9, 10, 11]. These models postulate that the present-day DM density arises due to an asymmetry in the DM particle-antiparticle density, similar to the baryon asymmetry, and that these were generated by similar mechanisms. This microphysical model is contrasted with that of the WIMP dark matter [12], in which the similarity in dark and visible matter densities must be taken as a coincidence [9, 13, 14]. Additionally, galactic DM distributions strongly suggest a DM candidate more complex than a collisionless WIMP. [15]
Certainly, ADM is not the only approach that invokes a common origin for baryonic and dark matter; instead, there exists the possibility that DM is not a separate fundamental field, but is instead a new, exotic state of visible, Standard Model matter. Such a proposal has precedent, with previous ideas ranging from quark nuggets, to strange matter (known as strangelets) [16, 17] to QCD balls [18]. Under the influence of an unspecified interaction, SM fermions may pair to form condensates that behave as scalar fields, analogously to the Cooper pairing [19] of BCS theory [20, 21].
A priori DM condensate candidates made of SM fermions and sustained by SM interactions necessarily involve neutrinos or quarks (see discussion in section II). However, since neutrinos were in thermal equilibrium at early times and space-filling quark condensates imply electrically charged DM, we are forced to introduce extra fermionic species to get a condensate that satisfy cosmological constraints. We shall assume that these species are produced in the early universe, preferentially out of equilibrium, but more generically with a large chemical potential-to-temperature ratio. While there are proposals in which the condensate is accomplished via a dark gauge interaction [4, 22], in this paper, the fermion interacts via SM boson mediation.
The paper is organized as follows. In the next section, we investigate possible attractive channels for the condensation within the SM. In section III, we discuss the effective theory for the Higgs mode of the condensate and translate cosmological constraints on DM to constraints on its parameters. In section IV, we specify the fermions giving rise to the condensate and in section V we conclude with a discussion on future directions. Throughout the paper, we work with the mostly plus metric signature convention and assume natural units when otherwise not mentioned.
II Attractive channel from Standard Model dynamics and fermion candidates
It is instructive to quickly explore all possible attractive channels of fermions in the Standard Model and understand why they cannot be a viable dark matter condensate. First, electrically charged leptons would give rise to electrically charged Cooper pairs, so we do not consider them (moreover they would modify the electroweak vacuum which is tightly constrained by EW precision measurements). Therefore, with SM fermions, we have two possibilities: quarks and neutrinos. We comment on known results on condensates of these fermions in the following sections.
SM quark condensate:
Color superconductivity [23, 24, 25, 26, 27, 28] is a well-studied phenomenon in the context of QCD at high densities. Quarks are in the fundamental of $\text{SU}(3)_{\text{c}}$ gauge group, so a $\langle qq\rangle$ condensate could necessarily be in a color antitriplet or a sextuplet because the tensor product of the color group fundamental representations splits as $\mathbf{3}\otimes\mathbf{3}=\bar{\mathbf{3}}\oplus\mathbf{6}$ [29]. Due to asymptotic freedom [30, 31], at high densities we can consider QCD at weak coupling [32], and single gluon exchanges generate an attractive potential between quarks in the $\bar{\mathbf{3}}$ channel (see [33, 34] and the discussion below). However, the specific combination of quarks in the pair will also depend on the flavor. From the phase diagram of QCD with three flavors [33], we expect that as we increase the chemical potential $\mu$ (from low enough temperatures), first there is a two-flavor superconducting (2SC) phase followed by a 3-color-flavor locking (CFL) superconducting phase as $\mu$ gets larger than the strange quark mass [33, 34]. However, even without increasing the chemical potential further, we already encounter a problem: the gap in those phases will be charged under the electroweak gauge group and its excitations will carry electric charges [35] (remarkably, in the CFL where chiral symmetry is broken, the number of Goldstone modes are the same as in the chiral vacuum at $\mu=0$ albeit with different masses and electric charges [36]). If the superconductor state is space-filling [^1] There are proposals [18, Ouyed:2023hqe] that assume the state to be confined in a spatial region (“quark clusters”). In this approach, the cluster total charge today reflects the mean charge density of the cosmic fluid at the time of its formation. So, quark clusters might have an overall small charge., DM would have a (large) electric charge and thus contribute to the electroweak vacuum, changing the $W$ and $Z$ masses and messing up the electroweak precision measurements. Thus, even without analyzing whether such pairings can happen in the early universe, there are difficulties for a quark superconductor to play the role of dark matter. Moreover, at finite temperature, we need a non-vanishing density of quarks to have a gap on the Fermi surface which would be a phenomenological catastrophe.
SM neutrino condensate:
At first sight, Standard Model neutrinos are promising candidates since they are electrically neutral and do not couple with gluons. Hence, $\langle\nu\nu\rangle$ would imply that the gap only couples with the $W$ and $Z$ bosons. The challenge here would be to find an attractive channel in the $\nu-\nu$ SM interactions and to explain why $\mu/T$ is large enough to allow for the superconductor phase. Such a channel has been previously explored [38, 39] but the thermal distribution of SM neutrinos coupled with their low chemical potential precludes the formation of any condensate at early times.
Since we wish to keep the SM bosonic content intact, let us discuss what attractive channels can be obtained with the SM interactions alone. As we mentioned, in color superconductivity the main attractive interaction that gives origin to the BCS instability is the single-gluon exchange between quarks. For quarks close to the Fermi surface at large chemical potentials, this interaction is described by weakly coupled Yang-Mills gauge theory [32]. Motivated by this, in this section, we want to find attractive channels among different fermionic states in the Standard Model. To do so, we recall a generic fermion-fermion interaction via a gauge boson exchange in an SU$(N)$ gauge theory with Dirac fermions in the fundamental of the gauge group. The amplitude for $\psi_{i}(p_{1})+\psi_{k}(p_{3})\to\psi_{j}(p_{2})+\psi_{\ell}(p_{4})$ is given by
$i\mathcal{M}$ $=(-g)\bar{u}_{j}(p_{2})\gamma^{\mu}T^{a}_{ji}u_{i}(p_{1})$
×$\times\frac{-i\left(\eta_{\mu\nu}+\frac{(p_{1}-p_{2})_{\mu}(p_{1}-p_{2})_{\nu}}{M^{2}}\right)}{(p_{1}-p_{2})^{2}-M^{2}}(-g)\bar{u}_{\ell}(p_{4})\gamma^{\nu}T^{a}_{\ell k}u_{k}(p_{3}),$ (1)
[…]
$T^{a}_{ji}T^{a}_{\ell k}$ $=T^{a}_{j(i}T^{a}_{|\ell|k)}+T^{a}_{j[i}T^{a}_{|\ell|k]}$
$=-\frac{N+1}{4N}\left(\delta_{ji}\delta_{k\ell}-\delta_{jk}\delta_{i\ell}\right)+\frac{N-1}{4N}\left(\delta_{jk}\delta_{i\ell}+\delta_{ji}\delta_{k\ell}\right),$ (2)
and so the color-antisymmetric channel $[ik]$ is attractive [33, 34]. If we had many fermionic flavors, then the anticommutative nature of the fermions would also force the attractive channel to be antisymmetric in the flavor space as well. One can also have a Higgs-Yukawa interaction between the vector-like fermions, which includes attractive chiral and non-chiral channels [39, 40, 41].
From the discussion above, we find that the $\nu$$\nu$ interaction via $Z$ boson exchange is repulsive. Neutrino condensation via Higgs attractive exchange has been explored in [39] after assuming a right-handed component, and in [38] it was shown that there is no attractive channel for neutrino pairing via SM gauge boson exchanges, while the $Z$ exchange may give rise to a neutron-neutrino condensate in [42]. From the previous discussion about the quark sector, we conclude that there cannot be an electrically neutral bound state made entirely of SM fermions paired via single $\text{SU}(3)_{\rm c}\times\text{SU}(2)_{\rm L}\times\text{U}(1)_{Y}$-gauge boson interactions. We can instead consider adding extra fermions to the SM but keeping its gauge structure intact. The following two well-motivated dark matter candidates that are minimal extensions will be considered in this work:
Vector-like quark condensate:
A minimal non-chiral extension would be to consider vector-like quarks (VLQs) neutral under the electroweak sector. Vector-like quarks, so named because they are left-right symmetric with respect to the gauge group, as opposed to the chiral structure of the SM quarks, are a well-studied and motivated anomaly-free extension to the matter content of the Standard Model (see [43] and references therein). VLQs have the benefit that their vector-like structure allows for the existence of a VLQ mass scale which can be set independently from the electroweak scale. Normal treatments usually set $v_{\rm EW}<m_{\rm VLQ}\sim\mathcal{O}(1\text{ TeV})$ to explain the non-observation of VLQs at the LHC to date [44, 45]. This freedom can be exploited in other ways, perhaps by setting it to a small value that allows the VLQs to condense under single-gluon exchange. The Higgs excitations of such a condensate could then play the role of a DM candidate.
Higgs-Induced Right-handed Neutrino condensate:
Right-handed neutrinos are a natural extension to the Standard Model, not least because their inclusion would restore the symmetry between quarks and leptons [46, 47]. The inclusion of these fermions would enable a Yukawa coupling between the left- and right-handed neutrinos, which provides the necessary attractive channel. Furthermore, since right-handed neutrinos are sterile relative to the rest of the SM, their cosmological evolution could decouple from radiation and they would remain in a non-thermal distribution that may preserve an appreciable chemical potential.
III Dark matter from superconductor states
Here we discuss how DM can be identified as the Higgs mode of the condensate. We start from a generic effective theory for the interacting fermions and discuss the particular quark and neutrino scenarios in sections (IV.1) and (IV.2). The starting point is a Nambu-Jonas-Lasinio-like fermion Lagrangian [48, 49] with an attractive four-fermion interaction of the form which is induced from an attractive channel of fermions.
$\mathcal{L}=-\bar{\psi}\left(\not{\partial}-m_{\psi}-\mu\gamma^{0}\right)\psi+\frac{g^{2}}{M^{2}}\left(\bar{\psi}\Gamma^{A}\psi\right)\left(\bar{\psi}\Gamma_{A}\psi\right).$ (3)
Introducing Hubbard-Stratonovich fields $\Phi_{\alpha}$ [50, 51], we can write the theory above as one without a four-fermion interaction
$\mathcal{L}_{\Phi}$ $=-\bar{\psi}\left(\not{\partial}-m_{\psi}-\mu\gamma^{0}\right)\psi$
$+\Phi_{\alpha}\left(\psi^{\dagger}\mathcal{O}^{\alpha}\psi^{*}\right)+\Phi_{\alpha}^{*}\left(\psi^{\text{T}}{\mathcal{O}^{\alpha}}^{*}\psi\right)-\frac{M^{2}}{g^{2}}\Phi_{\alpha}{\Phi^{\alpha}}^{*}.$ (4)
The action corresponding to $\mathcal{L}$ can be recovered after using the equations of motion for $\Phi_{\alpha}$ and $\Phi^{*}_{\alpha}$. Since we wish to apply the following results for different attractive channels in different theories, we shall write the formal dependence on $\mathcal{O}_{\alpha}$ and the index $\alpha$ explicitly.
In the Nambu-Gorkov basis [52, 53]
$\Psi=\begin{pmatrix}\psi\\ \psi^{*}\end{pmatrix},$ (5)
we have
$\mathcal{L}_{\Phi}=-\frac{1}{2}\bar{\Psi}S^{-1}_{F}(\Phi_{\alpha},m_{\psi},\mu)\Psi-\frac{M^{2}}{g^{2}}\Phi_{\alpha}{\Phi^{\alpha}}^{*},$ (6)
where
$S_{F}^{-1}(\Phi_{\alpha},m_{\psi},\mu)$
$=\begin{pmatrix}\gamma^{\mu}\partial_{\mu}-m_{\psi}-\mu\gamma^{0}&\Phi_{\alpha}i\gamma^{0}\mathcal{O}^{\alpha}\\ \Phi_{\alpha}^{*}i\gamma^{0}{\mathcal{O}^{\alpha}}^{*}&i\gamma^{0}{\gamma^{\mu}}^{\text{T}}i\gamma^{0}\partial_{\mu}-m_{\psi}-\mu\gamma^{0}\end{pmatrix}$ (7)
In the mean-field approximation, an effective action for $\Phi$ is obtained after integrating out the fermionic fields,
$e^{iS_{\text{eff}}(\Phi_{\alpha},\Phi_{\alpha}^{*})}$
$=\int\left[\mathrm{d}\bar{\Psi}\mathrm{d}\Psi\right]e^{-i\int\mathrm{d}^{4}x\left(\frac{1}{2}\bar{\Psi}S^{-1}_{F}(\Phi_{\alpha},m_{\psi},\mu)\Psi+\frac{M^{2}}{g^{2}}\Phi_{\alpha}{\Phi^{\alpha}}^{*}\right)}$ (8)
and the gap equation is
$\frac{\delta S_{\text{eff}}}{\delta\Phi_{\alpha}}=0,$ (9)
and its solution fixes the value of $\Phi_{\alpha}=\bar{\Phi}_{\alpha}(\mu,m_{\psi},M^{2}/g^{2})$ in terms of the parameters of the original Lagrangian. The gap $\bar{\Phi}_{\alpha}$ determines the value of the condensate:
$\langle\psi^{\text{T}}\mathcal{O}^{\alpha}\psi\rangle\propto\bar{\Phi}^{\alpha}.$ (10)
One core fact of BCS theory is that there is always a solution to the gap equation whenever the interaction term corresponds to an attractive channel [54, 55]. Moreover, any repulsive channel becomes irrelevant to the computation of the gap. This can be more easily seen from a Wilsonian perspective, in which modes away from the Fermi surface are integrated out, and repulsive interaction terms become irrelevant at those energies [54, 55].
In this paper, we propose to identify dark matter as the Higgs mode $\rho_{\alpha}(x^{\mu})$ of the gap. To get the microscopic DM parameters, such as mass and self-interaction couplings, we will then need an effective theory for fluctuations around $\bar{\Phi}_{\alpha}$:
$\Phi_{\alpha}(x^{\mu})=\left[\bar{\Phi}_{\alpha}+\rho_{\alpha}(x^{\mu})\right]e^{i\phi(x^{\mu})},$ (11)
where $\phi$ is the Goldstone mode of the condensate. This mode will be associated to a symmetry breaking by (10), which varies for the neutrino and vector-like quark scenarios. The effective field theory for the fluctuation modes was studied in [56] by perturbing around the ground state of the resulting effective action in (III). In terms of the canonically normalized field $\varphi=\left(\sqrt{3}/2\pi\right)\left(\mu/\Phi\right)\rho$, the Lagrangian for the Higgs mode is then [56, 4]
$\mathcal{L}=-\frac{1}{2}(\partial\varphi_{\alpha})^{2}-16\bar{\Phi}^{2}_{\alpha}\varphi_{\alpha}^{2}-\sum_{n\geq 2}\tilde{c}_{n}\frac{\bar{\Phi}_{\alpha}^{2}}{\mu^{n-2}}\varphi_{\alpha}^{n},$ (12)
where $\tilde{c}_{n}$ are order unity constants. From the Lagrangian above one identifies the DM mass $m$ and quartic coupling $\lambda$ as
$m=4\sqrt{2}\bar{\Phi},\quad\lambda\simeq\frac{\bar{\Phi}^{2}}{\mu^{2}}.$ (13)
In both scenarios explored in section IV, we introduce extra fermions on top of the SM leptons and quarks (however, the interaction giving rise to the four-fermion term in (3) still comes from the SM bosons). We shall assume these fermions were produced in the early universe, possibly during lepto- or baryogenesis. For generality, we forgo the details of production but rather comment on constraints on the temperature and chemical potential of the fermionic distribution. To get a Fermi surface, the fermions should be in the degenerate limit, with chemical potential much larger than the temperature, $\mu/T\gg 1$. Motivated by the Sakharov conditions [57] and electroweak baryogenesis [58, 59, 60, 61], these fermions may have been produced out of equilibrium and we can neglect the temperature corrections at leading order.
The energy density of the fermion fluid will be constrained by cosmological history. In the standard cosmic history context, adding an extra fluid before BBN corresponds to a change in the effective number of relativistic species, $N_{\text{eff}}$, which is set by the relativistic particle content of the SM. A shift $\Delta N_{\text{eff}}$ of this value is constrained by cosmological observations [8, 62]:
$\Delta N_{\text{eff}}<0.17\quad(\text{CMB}+\text{BAO})\;\text{at}\;95\%\;\text{CL}.$ (14)
We now use this constraint to find a maximum value for the Fermi energy of the degenerate gas. In the limit $\mu/T\gg 1$, the number and energy densities for a Fermi (ideal) gas are
$n_{\psi}$ $=\frac{g}{6\pi^{2}}\left(\mu^{2}-m_{\psi}^{2}\right)^{3/2},$ (15)
$\rho_{\psi}$ $=\frac{g}{16\pi^{2}}\left[\mu\left(2\mu^{2}-m_{\psi}^{2}\right)\sqrt{\mu^{2}-m_{\psi}^{2}}-m_{\psi}^{4}\text{arcosh}\left(\frac{\mu}{m_{\psi}}\right)\right],$ (16)
where $g$ is the number of degrees of freedom of each fermionic particle; each extra fermion flavor contributes a value of 2 to $g$. Note that for $\mu\gg m_{\psi}$, we have $n_{\psi}\propto\mu^{3}$ and $\rho_{\psi}\propto\mu^{4}$.
In an Friedmann-Lemaître-Robertson-Walker (FLRW) background, the chemical potential dependence of the scale factor can be obtained from particle number conservation. Using the conservation of the cosmic fluid’s entropy to write scale factor dependence in terms of temperature, we have (see [63] for details)
$\mu(T)=m_{\psi}\sqrt{1+\left(\frac{g_{*S}(T)}{g_{*S}(T_{\psi})}\right)^{2/3}\left(\frac{\mu(T_{\psi})^{2}-m_{\psi}^{2}}{T_{\psi}^{2}}\right)\left(\frac{T}{m_{\psi}}\right)^{2}},$ (17)
where $g_{*S}(T)$ is the effective number of degrees of freedom in entropy at temperature $T$ and $T_{\psi}$ is the temperature of the cosmic fluid when the non-thermal fermions were created. The evolution of the fermionic energy density is given by $\rho_{\psi}(T)=\rho_{\psi}(\mu(T))$. At early times, when $T\gg m_{\psi}$, $\rho_{\psi}$ redshifts as radiation,
$\rho_{\psi}(T)\simeq\frac{g}{8\pi^{2}}\left(\frac{\mu(T_{\psi})}{T_{\psi}}\right)^{4}\left(\frac{g_{*S}(T)}{g_{*S}(T_{\psi})}\right)^{4/3}T^{4},$ (18)
where we assumed $\mu(T_{\psi})\gg m_{\psi}$. After the temperature of the cosmic radiation drops below $m_{\psi}$, $\rho_{\psi}\propto T^{3}$ and we have a dust-like equation of state.
The shift in $N_{\text{eff}}$ during BBN is then [64]
$\Delta N_{\text{eff}}=\frac{8}{7}\frac{30g}{16\pi^{4}}\left(\frac{\mu(T_{\psi})}{T_{\psi}}\right)^{4}\left(\frac{g_{*S}(T_{\text{BBN}})}{g_{*S}(T_{\psi})}\right)^{4/3}.$ (19)
As long as $\Delta N_{\text{eff}}$ is less than the upper bound in (14), the mild cosmic history modification due to the presence of the fermion fluid satisfies the current cosmological constraints. Hence, (14) implies that $\mu(T_{\psi})$ should be at most of order $T_{\psi}$:
$\frac{\mu(T_{\psi})}{T_{\psi}}\lesssim\frac{1.22}{g^{1/4}}.$ (20)
The DM self-interaction cross-section $\sigma$ is constrained to satisfy $\sigma/m<0.1\;\text{cm}^{2}/\text{g}$ (see [65] and references therein). Using the expression for the mass and quartic coupling (13), we have
$\frac{m}{\text{eV}}\lesssim 10^{-19}\left(\frac{\mu_{0}}{\text{eV}}\right)^{4}$ (21)
where $\mu_{0}$ represents the Fermi energy today. There is a range of possible DM masses depending on the chemical potential right after the fermion production. From (17), we have
$\frac{m}{\text{eV}}$ $\lesssim 10^{-19}\left(\frac{m_{\psi}}{\text{eV}}\right)^{4}\times$
$\left[1+\left(\frac{g_{*S}(T_{0})}{g_{*S}(T_{\psi})}\right)^{2/3}\left(\frac{\mu(T_{\psi})^{2}-m_{\psi}^{2}}{T_{\psi}^{2}}\right)\left(\frac{T_{0}}{m_{\psi}}\right)^{2}\right]^{2}$ (22)
Using the constraint (20) with $\mu(T_{\psi})\gg m_{\psi}$, we get
$\frac{m}{\text{eV}}\lesssim 10^{-19}\left(\frac{m_{\psi}}{\text{eV}}\right)^{4}\left[1+\frac{1.22}{g^{1/4}}\left(\frac{g_{*S}(T_{0})}{g_{*S}(T_{\psi})}\right)^{2/3}\left(\frac{T_{0}}{m_{\psi}}\right)^{2}\right]^{2},$ (23)
which for $T_{0}\ll m_{\psi}$ gives
$\frac{m}{\text{eV}}\lesssim 10^{-19}\left(\frac{m_{\psi}}{\text{eV}}\right)^{4}.$ (24)
Hence, for a wide range of fermion masses, this model satisfies the self-interaction constraints provided DM is ultralight [66, 67] and consequently the DM phenomenology resembles the one for ultralight axions (see [68, 69, 70] and references therein). The neutrino scenario has a smaller mass, $m_{\psi}\sim 0.1\;\text{eV}$, compared to the vector-like quark case in which $m_{\psi}>10^{13}\;\text{eV}$ [44, 45]. However, from the specific form of the gap for each case, the DM mass can take on a wider range of values, corresponding to anything from ultralight DM to WIMPs [71] (and references therein), all depending primarily on the chemical potential (for the neutrino case) and the gauge coupling (for the VLQs case), as we see in section IV. The ultralight DM region is phenomenologically interesting because it would reproduce many aspects of superfluid dark matter (see e.g. [72] and references therein).
We can estimate which values of $\mu_{\psi}$ can explain the coincidence $\Omega_{c}\sim 5\Omega_{b}$ after assuming that the number density of fermions right after they are created $n_{\psi}(T_{\psi})$ is of the same order of the net baryon number density $n_{b}=\eta n_{\gamma}$ at that time, where $\eta$ is the baryon-to-photon ratio and $n_{\gamma}$ is the photon number density. The present baryonic density parameter is related to $\eta(T_{0})$ as
$\Omega_{b}h^{2}\simeq 3.66\times 10^{7}\eta,$ (25)
and BBN an CMB data implies $\eta(T_{0})\simeq 6\times 10^{-10}$ [8, 62, 73]. So, the baryonic density can be explained as long as we reproduce the observed value for $\eta(T_{0})$. From
$n_{b}(T_{\psi})=\xi_{\psi}(T_{\psi})n_{\psi}(T_{\psi}),$ (26)
and the expression for the fermionic number density in (16), we get
$\left(\frac{\mu_{\psi}^{2}-m_{\psi}^{2}}{T^{2}_{\psi}}\right)^{3/2}=\frac{12\zeta(3)}{g}\frac{\eta(T_{\psi})}{\xi_{\psi}}.$ (27)
From
$\eta=\frac{n_{\text{B}}}{n_{\gamma}}=\frac{\pi^{4}}{45\zeta(3)}g_{*S}(T)\frac{n_{\text{B}}}{s},$ (28)
where $s$ is the radiation entropy density, we have
$\left(\frac{\mu_{\psi}^{2}-m_{\psi}^{2}}{T^{2}_{\psi}}\right)^{3/2}=\frac{12\zeta(3)}{g}\frac{g_{*S}(T_{\psi})}{g_{*S}(T_{0})}\frac{\eta(T_{0})}{\xi_{\psi}}.$ (29)
For large $\mu_{\psi}$ and $g_{*S}(T_{\psi})\sim\mathcal{O}(10^{2})$,
$\frac{\mu_{\psi}}{T_{\psi}}\simeq\frac{1}{(g\xi_{\psi})^{1/3}}6\times 10^{-3}.$ (30)
As $\xi_{\psi},\,g\sim\mathcal{O}(1)$ we conclude that the relation between the baryonic and cold matter densities is reproduced provided $\mu_{\psi}$ is around 3 orders of magnitude smaller than the cosmic radiation temperature when the fermions are created. Note that (30) and (20) are compatible; hence, under the assumptions of this section, a degenerate gas of fermions in the early universe can indeed explain the $\Omega_{c}\sim 5\Omega_{b}$ coincidence today.
In the following subsections, we will find a wide range of parameters compatible with the above constraints, generating a broad spectrum of DM masses. In general, we expect phenomenology akin to the superfluid DM (see e.g. [74] and references therein); namely, we expect the large-scale evolution of the DM background energy density to be described by the fermionic one, while the scalar nature of the Higgs modes will play a leading role on galactic and sub-galactic scales, evading constraints on fermionic DM candidates, such as the Tremaine-Gunn bound [75].
We propose two independent scenarios that realize dark matter from fermionic condensates: right-handed sterile neutrinos and vector-like quarks [^2] By “sterile” we mean that the right-handed neutrinos are singlets of the SM gauge groups. Strictly speaking, in the neutrino case we have a superfluid condensate, but our formalism generally accommodates superconductor states. For instance, in the vector-like quark scenario we have a superconductor condensate, since these quarks are charged under the SM color group.. These are natural extensions of the SM fermion sector, with no changes in its bosonic sector.
Since the particulars, from the form of the gap to the allowed chemical potentials and associated densities, are different for the neutrino and vector-like quark scenarios, we will comment on constraints on the consistency of the values for $\Delta$ and $m$ individually for each fermion case covered in the following section.
IV Possible Scenarios
IV.1 Right-handed neutrino condensate
The notion of a neutrino condensate mediated by a scalar field between left- and right-handed neutrinos is not a new idea. However, most treatments of this idea miss an important subtlety which opens the possible parameter space. Yukawa interactions generically couple left- and right-handed fermion fields; this would in turn lead to the idea that any resulting condensate should consist of left-right (LR) bound states. However, it has been shown that the Yukawa interaction can mediate the formation of a condensate in both the $t$ channel [39] (the expected LR condensate) and the $s$ channel [40, 41] (leading to the possible formation of LL and RR condensates).
Given that the replusive $Z$ exchange is irrevelant at finite density, the relevant Lagrangian for the neutrino-Higgs sector, written in terms of Weyl spinors, is
[…]
where $\nu$ is the left-handed (Standard Model) neutrino and $\bar{\nu}$ is the right-handed neutrino. Upon inserting the Higgs vev and integrating out the resulting interaction, we obtain the effective Lagrangian for Dirac neutrinos at finite densities and below the Higgs mass scale,
$\mathcal{L}=-\bar{\psi}\left(\not{\partial}-m-\mu\gamma^{0}\right)\psi+\frac{y_{\nu}^{2}}{M_{h}^{2}}\left(\bar{\psi}\psi\right)\left(\bar{\psi}\psi\right),$ (32)
where $\psi^{\rm T}=(\nu,\bar{\nu}^{\dagger})$ is the Dirac spinor containing the neutrino fields, $y_{\nu}$ is the neutrino Yukawa coupling and $M_{h}$ is the Higgs mass. We will see that condensates containing both the same and different chiralities are possible using this effective Lagrangian. The resulting gap may be represented diagrammatically as in Figure 1.
Figure 1: Diagrammatic representation of a LL or RR gap arising from the Yukawa interaction between left- and right-handed neutrinos. The Higgs (dashed line) induces an attractive four-fermi interaction. The grey blob denotes a sum over neutrino ladder diagrams.
This action is equivalent to (3) with one quartic-coupling operator $\Gamma^{1}=\mathbb{1}$ and $g^{2}/M^{2}=y_{\nu}^{2}/M_{h}^{2}$. In the appendix A, we use a Fierz rearrangement to show that the only operators $\mathcal{O}_{\alpha}$ corresponding to attractive channels are $\gamma^{2}\gamma_{5}/4$ and $\gamma^{1}\gamma^{3}/4$, such that there are two possible gaps:
$\langle\psi^{\dagger}\gamma^{2}\gamma_{5}\psi\rangle\propto\Phi_{\rm LR},\quad\langle\psi^{\dagger}\gamma^{1}\gamma^{3}\psi\rangle\propto\Phi_{\rm LL+RR}.$ (33)
[…]
$\Phi_{\rm RR}=2\sqrt{M_{h}(\mu-m_{\psi})}\;e^{-\frac{\pi^{2}}{2\kappa_{R}^{2}\mu\sqrt{\mu^{2}-m_{\psi}^{2}}}},$ (34)
where $\kappa_{R}^{2}\simeq y_{\nu}^{2}/M^{2}_{h}$. The exponential dependence on the inverse coupling squared $\Phi\propto e^{-1/y_{\nu}^{2}}$ is expected from BCS theory. Note that the condensate breaks the lepton number symmetry spontaneously.
For tuned values of the chemical potential at early times, the condensate’s Higgs mode reproduces the dark nature of DM. For $m_{\psi}\approx 0.1\,\text{eV}$ and $\mu_{\psi}\approx 1.5\times 10^{14}\;\text{GeV}$, we have $\Phi_{\rm RR}\approx 10^{-20}\;\text{eV}$, such that the bound (III) is compatible with (13). At later times, the cosmological bounds will continue to be satisfied, because as $\mu$ gets closer to $m_{\psi}$, $\Phi_{\rm RR}$ is further suppressed. Note that the gap depends mainly on the chemical potential value. The matter of generating such a chemical potential is an open question in this work, but possible means of addressing this question have been examined in e.g. [77], wherein the inflaton’s VEV can act as a fermionic chemical potential.
IV.2 Vector-Like Quark condensate
The vector-like quark scenario would involve the inclusion of new fermion fields charged under $SU(3)_{\rm c}$ but which are otherwise singlets with respect to the Standard Model. If we consider a scenario with three VLQ flavors, then for a sufficiently high chemical potential $\mu\gg m_{\rm VLQ}$ we should expect the formation of a 3-color-flavor locking superconducting phase, mediated by a single gluon exchange.
Consider the Lagrangian for VLQs at finite densities,
[…]
$\qquad\qquad+\frac{G}{2}\left(\bar{\psi}\gamma_{\mu}T_{a}\psi\right)\left(\bar{\psi}\gamma^{\mu}T_{a}\psi\right).$ (35)
This action is equivalent to is of the form (3) with one quartic-coupling operator $\Gamma^{1}=\gamma_{\mu}T_{a}$ and $g^{2}/M^{2}=\frac{G}{2}$. The color piece can be extracted through factorization into a symmetric and antisymmetric piece; the remainder can be reworked using the Fierz identities, yielding the operator $\mathcal{O}_{\alpha}\propto\gamma_{0}\gamma_{2}$ corresponding to the dominant attractive channel [78, 79, 33]. The ground state of this theory contains a condensate of the form
$\Braket{\psi^{a\alpha}_{i{\rm L,R}}(\vec{p})\psi^{b\beta}_{j{\rm L,R}}(-\vec{p})\epsilon_{ab}}=\pm\Delta(p^{2})\epsilon^{\alpha\beta A}\epsilon_{ijA},$ (36)
where $\psi^{\alpha}_{i{\rm L,R}}$ denotes two-component Weyl spinors with flavor and color indices $i$ and $\alpha$, respectively. It has been shown that a condensate of the CFL form is the dominant ground state in three-flavor QCD at finite densities [36, 80, 81, 82, 83]. Under this scenario, the VLQs have a microscopic symmetry of the form $G=\text{SU}(3)_{\rm L}\times\text{SU}(3)_{\rm R}\times\text{U}(1)_{B}$; in the CFL phase this is broken to $G^{\prime}=\text{SU}(3)_{\rm c+L+R}\times\mathbb{Z}_{2}$. As with the ordinary quarks, the chiral flavor symmetry is also broken, except here it is ‘locked’ to the color symmetry, necessary due to the pairing of quarks of like helicity. Such a symmetry breaking does induce a small gluon mass. From a phenomenological perspective, non-zero gluon masses of up to an MeV are consistent with experimental measurements [84]. Before proceeding further, it is worth examining whether such a condensate can form.
The dominant interaction in this model will be of single-gluon exchange within the VLQ sector (contrast with the neutrino case where the condensate forms in the RH neutrino sector, via a Higgs interaction with the LH neutrinos). Single-gluon exchange diagrams lead to two different interactions, one attractive (forming the condensate) and one repulsive (thermalizing the quarks). The ratio of these interactions is given by $\mathcal{M}_{a}/\mathcal{M}_{s}=(N_{c}+1)/(N_{1}-1)$ (from (II)). For $N_{c}=3$ (as we are working with standard QCD), then the interaction rates of the two processes $\Gamma=n\sigma v$ gives the ratio of time-scales $t_{a}/t_{s}=1/4$. Thus the time-scale for condensate formation is shorter than that for scattering.
Now we turn to the question of whether the condensate can form cosmologically. In order for the interaction to be appreciable within a Hubble patch, we expect $\Gamma/H\gg 1$. For gluon exchange dominated by modes near the Fermi surface [4]
$\frac{\Gamma}{H}=\frac{n_{\psi}\alpha^{2}}{k^{2}_{F(\psi)}H_{\psi}}\left(\frac{a(t)}{a(t_{\psi})}\right)$ (37)
where as before the $\psi$ subscript corresponds to the quantity evaluated at the time of the VLQ production, and $\alpha=g^{2}/4\pi$ is the QCD coupling. If we assume most fluctuations have $k=k_{F}$, then we can estimate $\rho_{c}\approx nk_{F}$. Furthermore, we can write an upper bound on $k_{F}$ to be $H_{\psi}$. Lastly, using the Friedmann equation we can write $H_{\psi}=\sqrt{\rho_{{\rm tot},\psi}}/(\sqrt{3}M_{\rm Pl})$, where $\rho_{{\rm tot},\psi}$ is the energy density of the universe at the time of VLQ production. Putting these together we can write
$\frac{\Gamma}{H}\sim\frac{\rho_{c}}{\rho_{{\rm tot},\psi}}\frac{M_{\rm Pl}^{2}}{H_{\psi}^{2}}\left(\frac{a(t)}{a(t_{\psi})}\right)$ (38)
Since the condensate forms deep in radiation-domination, we can write $\rho_{c}=\rho_{\psi}\sim g\mu(T_{\psi})^{4}/8\pi^{2}$ using (18) while $\rho_{{\rm tot},\psi}\sim\pi^{2}g_{*}(T_{\psi})T_{\psi}^{4}/30$, where $g_{*}(T)$ is the number of degrees of freedom in radiation at temperature $T$. Thus we arrive at
$\frac{\Gamma}{H}$ $\sim\frac{15}{4\pi^{4}}\frac{g}{g_{*}(T_{\psi})}\left(\frac{\mu(T_{\psi})}{T_{\psi}}\right)^{4}\frac{M_{\rm Pl}^{2}}{H_{\psi}^{2}}\left(\frac{a(t)}{a(t_{\psi})}\right)$
$\sim 10^{84}\left(\frac{a(t)}{a(t_{\psi})}\right)\left(\frac{\mu(T_{\psi})}{T_{\psi}}\right)^{4},$ (39)
where we used $g_{*}(T_{\psi})\sim\mathcal{O}(10^{2})$ so that $g/g_{*}(T_{\psi})\sim\mathcal{O}(10^{-2})$ and let $H_{\psi}\simeq H_{\rm BBN}\sim 10^{-44}M_{\rm Pl}$. Hence, due to the large numerical factor in the rightmost-hand side, the condition $\Gamma/H\gg 1$ is easily satisfied.
With the question of formation addressed, as found in [85] and discussed further in [33], the CFL phase gap is given by
$\Phi\sim b\frac{\mu}{g^{5}}e^{-\frac{3\pi^{2}}{\sqrt{2}g}}$ (40)
where the prefactor $b$ has been computed in the CFL phase to be $b\sim 512\pi^{4}2^{-\frac{1}{3}}\left(\tfrac{2}{3}\right)^{\frac{5}{2}}$ [81, 86].
Compared with the neutrino case, the greater lower bound on the VLQ mass $m_{\rm VLQ}>10^{13}\text{ eV}$ and the gap’s independence of the mediator mass gives greater leeway in this scenario to satisfy the DM constraints computed in section III. For $\mu\sim 10^{14}$ eV and $g\sim 0.23$ we find $\Delta\sim 10^{-20}$ eV, which is compatible with (24), but other mass ranges are possible depending mainly on the precise value of $g$.
V Discussion and Conclusion
In this work we have presented an analysis regarding whether dark matter states are viable within the structure of the Standard Model. Although it is not possible, due to precision data, to construct dark states using only Standard Model fields, it is possible to do so within the gauge structure of the Standard Model, by augmenting the matter sector with well-motivated additional fermionic fields, namely the sterile neutrino and the vector-like quark. Each scenario has its unique aspects, but much of the physics common to both scenarios arises from the well-known physics of superconductivity.
Postulating a degenerate fermionic background, for sufficient values of chemical potential, the formation of a condensate is inevitable given an attractive interaction. This presents the possibility that what we have heretofore called dark matter is simply a particular state of the fermionic species, mediated by the known interactions of the Standard Model. In this case, the dark matter particle would be the Higgs mode of the resulting superconducting state, similar to how the Higgs boson detected at the LHC is but the excitation of the background Higgs field.
The necessary chemical potential in the neutrino case is adjusted to account for the fact that the neutrino mass is bounded from above and that the interaction is mediated by the Higgs which is relatively heavy for such an interaction. Due to the constraints on the neutrino mass and the Higgs mass, the possible values for the DM particle are fixed primarily by the chemical potential and it is thus possible in this case to generate DM masses in the ultralight range. In the vector-like quark case, because of the higher possible VLQ mass values and the massless (or nearly so) nature of the mediating particle, the scenario is not as strongly tuned and can naturally accommodate a wide range of DM masses from the ultralight to WIMP-scale DM.
In our proposal, there is only one fluid, fundamentally made of fermions, but in a state that gives rise to different modes (Goldstone and Higgs excitations). The Goldstone mode is massless and so it redshifts as radiation right after the formation of the condensate, hence it is not relevant for the background energy density (similar to the model discussed in [4]). The DM particle consists of Higgs mode excitations of the background superconducting state. Since we have local energy conservation, the DM density can be obtained by redshifting the fermionic energy density (16) in the fermions right after their creation. Alternatively, we can calculate $\Omega_{c}$ at the time of condensate formation and compare it with Eq. (16). In that sense any bounds on $\Omega_{c}$ will translate into bounds on $\rho_{\psi}$. For instance, to get Eq. (30) we used the observed value of $\Omega_{b}$, its relation with $\Omega_{c}$, and the value of the baryon-to-photon ratio at early times.
It is worth commenting that the matter-radiation equality scale $T_{\rm eq}\sim 0.3\;\text{eV}$ is of the order of the cosmological upper bound on the sum of neutrino masses, $\sum_{i}m_{\nu,i}<0.13\;\text{eV}$ [8, 62]. For the neutrino case, $\rho_{\psi}$ starts to redshift as dust when $T\sim m_{\nu}$, i.e. just before $z_{\rm eq}$. Or, equivalently, neutrinos having an $\mathcal{O}(10^{-1})\;\text{eV}$ mass may be the reason why $z_{\rm eq}\sim\mathcal{O}(10^{3})$. Another possible line of investigation is whether the condensate can play any role in solutions to the Hubble tension (see [87, 88] for reviews), given the characteristic scale (eV) that new physics should have in order to solve the tension [89].
An important aspect left for future work is the phenomenology of this model, specifically examining such aspects as the core-cusp problem or the baryonic Tully-Fisher relation. For a wide range of parameters, we expect the small-scale phenomenology of the condensate to be similar to the superfluid DM case, thus avoiding usual astrophysical constraints on fermionic particle dark matter candidates [74]. Other phenomena associated with superfluid DM, such as vortices [90, 91, 92], are also possible and worth exploring in this context. It will be important to distinguish this model from a fundamental ultralight scalar dark matter candidate. One possibility for future work is to study the possible neutrino oscillations induced by the background neutrino condensate as neutrinos propagate through it. From a more theoretical perspective, one could also investigate how the four-fermion interaction affects the fermionic densities and get a fully-fledged expression for the condensate contribution to the dark matter density. Another direction would be explicitly constructing the baryogenesis model that gives origin to a degenerate fermion background with suitable chemical potential.
Acknowledgements.
We thank Chris Cappiello, Ian Dell’Antonio, Loukas Gouskas, Leah Jenks, Greg Landsberg, Evan McDonough, Philip Phillips, Michael Peskin, and David Spergel for useful discussions. We thank Jorge Sofo for inspiring SA for pursuing this project. HBG is supported by the National Science Foundation, MPS Ascending Fellowship, Award 2213126.
Appendix A Fierz rearrangement
We want to express the bilinear squared $(\bar{\psi}\psi)^{2}$ as a product of $\psi^{\dagger}\psi^{\dagger}$ and $\psi\psi$ to make particle-particle pairing in the mean field approximation. We have
$(\bar{\psi}\psi)^{2}=(\psi^{\dagger}i\gamma^{0}\psi)(\psi^{\dagger}i\gamma^{0}\psi)=\sum_{A,B}^{16}\xi_{AB}(\psi^{\dagger}\Theta^{A}\psi^{*})(\psi^{\text{T}}\bar{\Theta}^{B}\psi),$ (41)
where $\Theta^{A}$ are the elements of the Clifford algebra basis and only the six antisymmetric among those contribute to the summation. Typically, one works on the basis
$\Gamma^{A}$ =$=\{\mathbb{1},\gamma^{\mu},\gamma^{\mu\nu},\gamma^{\mu\nu\rho},\gamma^{\mu\nu\rho\sigma}\},$
$\Gamma_{A}$ =$=\{\mathbb{1},\gamma_{\mu},\gamma_{\nu\mu},\gamma_{\rho\nu\mu},\gamma_{\sigma\rho\nu\mu}\},\quad\sigma>\rho>\nu>\mu,$ (42)
where $\gamma^{\mu_{1}\cdots\mu_{r}}=\gamma^{[\mu_{1}}\dots\gamma^{\mu_{r}]}$. These matrices have normalization such that
$\text{tr}\left(\Gamma^{A}\Gamma_{B}\right)=4\delta^{A}_{B}.$ (43)
Note that $\{\Gamma^{A}\}$ is the basis for the vector times dual-vector representation. This can be seen from the fact that $\bar{\psi}\Gamma^{A}\chi$ is a scalar. On the other hand, $\Theta^{A}$ ($\bar{\Theta}^{B}$) denotes the basis elements in the vector-vector (dual-vector times dual-vector) representation, such that $\bar{\chi}\Theta^{A}\bar{\psi}^{\text{T}}=(\psi^{\text{T}}\bar{\Theta}^{A}\chi)^{\dagger}$ is a scalar. This relation also implies that $\bar{\Theta}^{A}=(i\gamma^{0})^{*}{\Theta^{A}}^{\dagger}(i\gamma^{0})^{*}$. The basis in both representations are related by $\Theta^{A}=\Gamma^{A}C$, where $C$ is the unitary charge conjugation matrix. Such a matrix enters in the symmetry property of $\Gamma^{A}$:
${\Gamma^{A}}^{\text{T}}=t_{0}t_{r}C\Gamma^{A}C^{-1},$ (44)
where $t_{r}=\pm 1$ depending on the rank of $\Gamma^{A}$. Choosing $t_{0}=1=-t_{1}$, we have $t_{2}=-1$ and $t_{3}=1=t_{4}$. In a representation in which ${\gamma^{\mu}}^{\dagger}=\gamma^{0}\gamma^{\mu}\gamma^{0}$ (hermitian representation) we have
${\gamma^{\mu}}^{*}=-t_{0}t_{1}B\gamma^{\mu}B^{-1},\quad B=it_{0}C\gamma^{0}.$ (45)
Using this result, it is straightforward to show that
${\Gamma^{A}}^{\dagger}=t_{1}t_{r}\gamma^{0}\Gamma^{A}\gamma^{0},$ (46)
and hence
$\bar{\Theta}^{A}=-t_{0}t_{r(A)}C^{\dagger}\Gamma^{A}.$ (47)
One can show that, with the choice of basis above, we have the normalization
$\text{tr}\left(\Theta^{A}\bar{\Theta}_{B}\right)=-t_{0}t_{r(B)}\text{tr}\left(\Gamma^{A}\Gamma_{B}\right)=-t_{0}t_{r(B)}4\delta^{A}_{B}.$ (48)
To get the typical particle-antiparticle Fierz rearrangement, we start with the identity
$\left(\Gamma^{C}\right)_{ij}\left(\Gamma^{D}\right)_{k\ell}=\sum_{A,B}C^{CD}_{AB}\left(\Gamma^{A}\right)_{i\ell}\left(\Gamma^{B}\right)_{kj}$ (49)
and multiply it by $\left(\Gamma_{E}\right)_{jk}\left(\Gamma_{F}\right)_{li}$ to find
$C^{CD}_{EF}=\frac{1}{16}\text{tr}\left(\Gamma^{C}\Gamma_{E}\Gamma^{D}\Gamma_{F}\right).$ (50)
Similarly, to get the particle-particle rearrangement, we start from
$\left(\Gamma^{A}\right)_{ij}\left(\Gamma^{B}\right)_{kl}=\sum_{C,D}D^{AB}_{CD}\left(\Theta^{C}\right)_{il}\left(\bar{\Theta}^{D}\right)_{kj}.$ (51)
To find
$D^{AB}_{FE}=\frac{1}{16}\text{tr}\left(\Gamma^{A}\Theta_{E}{\Gamma^{B}}^{\text{T}}{\bar{\Theta}}_{F}\right).$ (52)
So, from the relation between $\Theta$ and $\Gamma$, we have
$D^{AB}_{FE}$ $=-\frac{1}{16}t_{r(B)}t_{r(F)}\text{tr}\left(\Gamma^{A}\Gamma_{E}\Gamma^{B}\Gamma_{F}\right)$
$=-t_{r(B)}t_{r(F)}C^{AB}_{EF}.$ (53)
A glance at (41) reveals that $\xi_{AB}=-D^{00}_{AB}=-t_{r(F)}C^{00}_{BA}$. We now compute $C^{00}_{BA}$. Firstly, note that the trace of an odd number of $\gamma^{\mu}$ matrix vanishes and so will the trace of $\Gamma^{A}$’s with unmatched rank. Thus $C^{00}_{BA}\propto\delta_{BA}$. To compute the traces in (50), it is convenient to write the basis as $\Gamma^{A}=\{\mathbb{1},\gamma^{\mu},\gamma^{0i},\gamma^{ij},\gamma^{\mu}\gamma_{5},\gamma_{5}\}$ with $i<j$, where $\gamma_{5}=-i\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}$. Note that for $\Gamma_{A}$, the rank-2 elements have opposite index ordering compared to rank-2 $\Gamma^{A}$ elements, otherwise the elements will not be normalized as in (43). For instance,
$\left.\text{tr}(\gamma^{ij}\gamma_{k\ell})\right|_{\begin{subarray}{c}j>i\\ k>\ell\end{subarray}}=4\left(\eta^{ij}\eta_{k\ell}-\delta^{i}_{k}\delta^{j}_{\ell}+\delta^{i}_{\ell}\delta^{j}_{k}\right)_{\begin{subarray}{c}j>i\\ k>\ell\end{subarray}}=4\delta^{i}_{\ell}\delta^{j}_{k}$ (54)
would have an opposite sign if $\ell>k$. With that in mind, we collect the results for $D^{00}_{AB}$:
$D^{00}_{\mathbb{1}\mathbb{1}}$ $=t_{0}C^{00}_{\mathbb{1}\mathbb{1}}=-\frac{1}{4},$ (55)
$D^{00}_{00}$ $=t_{1}C^{00}_{00}=-\frac{1}{4},$ (56)
$D^{00}_{ij}$ $=t_{1}C^{00}_{ij}=-\frac{1}{4}\eta_{ij},$ (57)
$D^{00}_{i0\;j0}$ $=t_{2}C^{00}_{i0\;j0}=-\frac{1}{4}\eta_{ij},$ (58)
$D^{00}_{ij\;k\ell}$ $=t_{2}C^{00}_{ij\;kl}=-\frac{1}{4}\eta_{ik}\eta_{j\ell},$ (59)
[…]
$D^{00}_{i5\;j5}$ $=t_{3}C^{00}_{i5\;j5}=\frac{1}{4}\eta_{ij},$ (61)
$D^{00}_{55}$ $=t_{4}C^{00}_{55}=\frac{1}{4}.$ (62)
Using these results into (51), we find
$\left(\gamma^{0}\right)_{ij}\left(\gamma^{0}\right)_{k\ell}$ $=-\frac{1}{4}\left(C\right)_{ik}\left(-C^{\dagger}\right)_{\ell j}-\frac{1}{4}\left(\gamma^{0}C\right)_{ik}\left(C^{\dagger}\gamma^{0}\right)_{\ell j}$
$-\frac{1}{4}\eta_{mn}\left(\gamma^{m}C\right)_{ik}\left(C^{\dagger}\gamma^{n}\right)_{\ell j}$
$-\frac{1}{4}\eta_{mm}\left(\gamma^{0m}C\right)_{ik}\left(C^{\dagger}\gamma^{0n}\right)_{\ell j}$
$-\frac{1}{4}\eta_{mn}\eta_{pq}\left(\gamma^{mq}C\right)_{ik}\left(C^{\dagger}\gamma^{np}\right)_{\ell j}+$
$+\frac{1}{4}\left(\gamma^{0}\gamma_{5}C\right)_{ik}\left(-C^{\dagger}\gamma^{0}\gamma_{5}\right)_{\ell j}$
$+\frac{1}{4}\eta_{mn}\left(\gamma^{m}\gamma_{5}C\right)_{ik}\left(-C^{\dagger}\gamma^{n}\gamma_{5}\right)_{\ell j}+$
$+\frac{1}{4}\left(\gamma_{5}C\right)_{ik}\left(-C^{\dagger}\gamma_{5}\right)_{\ell j}.$ (63)
When we multiply this expansion by $\psi^{\dagger}_{i}\psi_{k}^{\dagger}\psi_{\ell}\psi_{j}$ to get (41), only the terms with antisymmetric $\Theta^{A}$ will contribute. The symmetry property of $\Theta^{A}$ depends on its rank:
${\Theta^{A}}^{\text{T}}=\left(\Gamma^{A}C\right)^{\text{T}}=t_{r}C^{2}\Gamma^{A}C=t_{r}C^{2}\Theta^{A}.$ (64)
We shall work concretely in the Weyl representation, where $C=-i\gamma^{0}\gamma^{2}$. In that case, $C^{2}=-\mathbb{1}$ and so $\Theta^{A}$ is antisymmetric for ranks such that $t_{r}=1$ which correspond to $r=\{0,3,4\}$. Then we are interested in four terms in the expansion above:
$\left(\gamma^{0}\right)_{ij}\left(\gamma^{0}\right)_{k\ell}$ $=-\frac{1}{4}\left(C\right)_{ik}\left(-C^{\dagger}\right)_{\ell j}$
$-\frac{1}{4}\left(\gamma^{0}\gamma_{5}C\right)_{ik}\left(-C^{\dagger}\gamma^{0}\gamma_{5}\right)_{\ell j}+$
$+\frac{1}{4}\eta_{mn}\left(\gamma^{m}\gamma_{5}C\right)_{ik}\left(-C^{\dagger}\gamma^{n}\gamma_{5}\right)_{\ell j}$
$+\frac{1}{4}\left(\gamma_{5}C\right)_{ik}\left(-C^{\dagger}\gamma_{5}\right)_{\ell j}+\cdots\;.$ (65)
Explicitly in terms of $\gamma$ matrices, we have
$\left(i\gamma^{0}\right)_{ij}\left(i\gamma^{0}\right)_{k\ell}$ $=-\frac{1}{4}\left(\gamma^{0}\gamma^{2}\right)_{ik}\left(\gamma^{0}\gamma^{2}\right)_{\ell j}$
$-\frac{1}{4}\left(\gamma^{2}\gamma_{5}C\right)_{ik}\left(\gamma^{2}\gamma_{5}\right)_{\ell j}-$
$-\frac{1}{4}\left(\gamma^{3}\right)_{ik}\left(\gamma^{3}\right)_{\ell j}-\frac{1}{4}\left(\gamma^{0}\gamma_{5}\right)_{ik}\left(\gamma^{0}\gamma_{5}\right)_{\ell j}$
$-\frac{1}{4}\left(\gamma^{1}\right)_{ik}\left(\gamma^{1}\right)_{\ell j}-$
$-\frac{1}{4}\left(\gamma^{1}\gamma^{3}\right)_{ik}\left(\gamma^{1}\gamma^{3}\right)_{\ell j}+\cdots\;.$ (66)
Finally, we express the particle-particle Fierz rearrangement in a manifestly real way as
$(\bar{\psi}\psi)^{2}$ $=\frac{1}{4}\left(\psi^{\dagger}\gamma^{0}\gamma^{2}\psi^{*}\right)\left[\psi^{\text{T}}\left(\gamma^{0}\gamma^{2}\right)^{*}\psi\right]$
$-\frac{1}{4}\left(\psi^{\dagger}\gamma^{2}\gamma_{5}\psi^{*}\right)\left[\psi^{\text{T}}\left(\gamma^{2}\gamma_{5}\right)^{*}\psi\right]+$
$+\frac{1}{4}\left(\psi^{\dagger}\gamma^{3}\psi\right)\left[\psi^{\text{T}}\left(\gamma^{3}\right)^{*}\psi\right]$
$+\frac{1}{4}\left(\psi^{\dagger}\gamma^{0}\gamma_{5}\psi\right)\left[\psi^{\text{T}}\left(\gamma^{0}\gamma_{5}\right)^{*}\psi\right]+$
$+\frac{1}{4}\left(\psi^{\dagger}\gamma^{1}\psi\right)\left[\psi^{\text{T}}\left(\gamma^{1}\right)^{*}\psi\right]$
$-\frac{1}{4}\left(\psi^{\dagger}\gamma^{1}\gamma^{3}\psi\right)\left[\psi^{\text{T}}\left(\gamma^{1}\gamma^{3}\right)^{*}\psi\right]$
$=\frac{1}{4}\sum_{a}\xi_{a}\left(\psi^{\dagger}O^{a}\psi\right)\left[\psi^{\text{T}}\left(O^{a}\right)^{*}\psi\right],$ (67)
with $\xi_{a}=1$ for $\{\gamma^{0}\gamma^{2},\,\gamma^{3},\,\gamma^{0}\gamma^{5},\,\gamma^{1}\}$ and $\xi_{a}=-1$ for $\gamma^{2}\gamma_{5}$ and $\gamma^{1}\gamma^{3}$.